OpenStudy (anonymous):

can someone help me with this problem 2/3-1/3+5/3 we got the answer but don't know how to show math for it

OpenStudy (mertsj):

Let's say you have 2 cats and one cat gets killed in the road so you adopt 5 more cats from the shelter. How many cats do you have now?

Directrix (directrix):

2/3-1/3+5/3. These are like terms. They all have the same last name of "third." 2/3 - 1/3+ 5/3 = (2 - 1 + 5) / 3 = 6/3 = 2.